# Orbit structures and complexity in Schubert and Richardson Varieties

### v1

@@ Abstract.
The goal of this paper is twofold. Firstly, we provide a type-uniform formula for the torus complexity of the usual torus action on a Richardson variety, by developing the notion of algebraic dimensions of Bruhat intervals, strengthening a type A result by Donten-Bury, Escobar and Portakal. In the process, we give an explicit description of the torus action on any Deodhar component as well as describe the root subgroups that comprise the component. Secondly, we exhibit a codimension preserving bijection between Levi-Borel subgroup (a Borel subgroup of a Levi subgroup in a reductive algebraic group) orbits in the big open cell of a Schubert variety and torus orbits in the big open cell of a distinguished Schubert subvariety. This bijection has many applications: a type-uniform formula for the Levi-Borel complexity of the usual Levi-Borel subgroup action on a Schubert variety as well as an indexing set for Levi-Borel subgroup orbits in a Schubert variety. We conclude by extending the Levi-Borel complexity results to a large class of Schubert varieties in the partial flag variety.
@@ 1.1. Group orbits in the full flag variety
Let $G$ be a complex, connected, reductive algebraic group of rank $r$. Fix a maximal torus $T$ in $G$, and a Borel subgroup $B$ of $G$ containing $T$. The homogeneous space $G/B$ is a smooth projective variety known as the full flag variety. The study of the flag variety first arose out of need to formalize and justify the enumerative geometry of H. Schubert as laid out in Hilbert’s 15th problem. These varieties are a central theme in much of the mathematics of the 20th century and beyond, with deep and fundamental connections to algebraic geometry, Lie theory, representation theory, algebraic combinatorics, and commutative algebra.
@@ 1.3. Levi-Borel subgroup orbits in Schubert Varieties
We now turn our attention to the orbits of Levi-Borel subgroups (defined below) in a Schubert variety. Our work uncovers a deep structural relationship between the Levi-Borel orbits in the big cell of a Schubert variety and the $T$-orbits in the big cell of a distinguished Schubert subvariety. Specifically, we exhibit a codimension-preserving bijection between these sets of orbits.
@@
Let $w\in W$, $I\subseteq\Delta$, and let $w={{}_{I}w}\prescript{I}{}{w}$ be the left parabolic decomposition of $w$ with respect to $I$. The map $\mathfrak{O}\!:\!\mathcal{O}_{T}(X^{\circ}_{\prescript{I}{}{w}})\!\rightarrow\!\mathcal{O}_{B_{L_{I}}}(X_{w}^{\circ})$ given by $\Theta\!\mapsto\!B_{L_{I}}\prescript{I}{}{w}x$, where $x$ is any point in $\Theta$, is a codimension-preserving bijection. That is,
@@
Theorem 1.4 allows us to intertwine the study of Levi-Borel orbits with the vast literature on torus orbits and varieties, including our own Corollary 1.3, and has myriad applications which we now detail. Our first application is Proposition 4.9 which provides a formula for the minimal codimension of a $B_{L_{I}}$-orbit in $X_{w}$. In general, the formula given in Proposition 4.9 is non-trivial to compute. However, in the case where the Levi subgroup $L_{I}$ acts on the Schubert variety $X_{w}$ the formula simplifies to a closed form.
@@
As a second application of Theorem 1.4, we are able to describe the orbits of Levi-Borel subgroups in the Schubert variety itself.
Corollary 1.6.
Let $w\in W$, $I\subseteq\Delta$, and let $w={{}_{I}w}\prescript{I}{}{w}$ be the left parabolic decomposition of $w$ with respect to $I$. There is a bijection between $\mathcal{O}_{B_{L_{I}}}(X_{w})$ and
$\bigsqcup_{u\leq w}\mathcal{O}_{T}(X^{\circ}_{\prescript{I}{}{u}})\times\{u\}$
We highlight that in the $I=\Delta$ case, $B_{L_{I}}=B$ and $\prescript{I}{}{u}=\mathrm{id}$ so each $\mathcal{O}_{T}(\prescript{I}{}{u})$ is a singleton set. Hence, in this case Corollary 1.6 recovers the Bruhat decomposition, namely that the Borel subgroup orbits in a Schubert variety are in bijection with the set of all $u\leq w$. In this way, Theorem 1.4 and Corollary 1.6 can be viewed as the first steps towards a framework that aims to generalize the study of Borel subgroup orbits in the full flag variety to the study of Levi-Borel subgroup orbits in Schubert varieties. Our next step, with work in progress, is understanding the inclusion order on these orbit closures. The combinatorial indexing set developed in this paper is an important cornerstone of this work. Throughout the years, the study of such Bruhat orders has led to much fruitful research. One related example, is that of the orbits of spherical Levi subgroups in Dynkin type $A$. These orbits are indexed by $(p,q)$-clans [MO90, Y97], and their inclusion order is described via clan combinatorics [W16].
@@
The outline of this paper is as follows. In Section 2, we introduce the necessary background, including Schubert geometry and Coxeter groups. In Section 3, we fully develop the notion of algebraic dimensions and establish Theorem 1.2. In Section 4, we prove Theorem 1.4, Theorem 1.5, and Corollary 1.6 with one of our main tools being an equivariant isomorphism that allows us to use results in Section 3. Section 4 concludes with an extension of the Borel-Levi subgroup complexity results to Schubert varieties in partial flag varieties.
@@ 2. Preliminaries
We first recall some preliminaries and set notation which will be useful throughout.
@@ 2.1. Combinatorics of the Weyl group
${\mathcal{I}}_{L}(w)=\{\alpha\in\Phi^{+}\>|\>w(\alpha)\in\Phi^{-}\},\quad{\mathcal{I}}_{R}(w)=\{\alpha\in\Phi^{+}\>|\>w^{-1}(\alpha)\in\Phi^{-}\},$
@@ 3. The torus complexity of Richardson varieties
In this section, we establish a formula for the complexity of the usual torus action on any Richardson variety $\mathcal{R}_{u,v}$ where $u\leq v$ in the Bruhat order.
@@ 4.1. An equivariant isomorphism
Set $U_{\bar{a}}=\prod_{\alpha\in\phi^{+}(I)\setminus{\mathcal{I}}(a)}U_{\alpha}$. Then $U_{\bar{a}}$ is a closed subgroup of $U$ normalized by $T$ and
$B_{L_{I}}=T\ltimes U_{L_{I}}=T\ltimes\prod_{\alpha\in\phi^{+}(I)}U_{\alpha}=T\ltimes(U_{a}U_{\bar{a}}).$
For a group $H$ acting on a variety $X$, denote the isotropy subgroup of $H$ at $x\in X$ by $H_{x}$.
@@
Let $x\in X_{w}^{\circ}$. Then $U_{\bar{a}}\leq B_{x}$.
@@
The big open cell $X_{w}^{\circ}$ is a $B$-orbit and so the isotropy subgroups $B_{y}$ are identical at every point in $y\in X_{w}^{\circ}$. Consider the isotropy subgroup at the point $wB\in X_{w}^{\circ}$. Then
$U_{\bar{a}}wB=ww^{-1}U_{\bar{a}}wB=w\prod_{\alpha\in\Phi^{+}(I)\setminus{\mathcal{I}}(a)}U_{w^{-1}(\alpha)}B,$
where the last equality is [J03, Part II, 1.4(5)]. Let $\alpha\in\Phi^{+}(I)\setminus{\mathcal{I}}(a)$. We claim $\alpha\notin{\mathcal{I}}(w)$. The claim would imply that $w^{-1}(\alpha)\in\Phi^{+}$, and continuing the above equation yields
$w\prod_{\alpha\in\Phi^{+}(I)\setminus{\mathcal{I}}(a)}U_{w^{-1}(\alpha)}B=wB,$
proving that $U_{\bar{a}}\leq B_{wB}=B_{x}$.
We now show ${\mathcal{I}}(w)\cap\Phi^{+}(I)={\mathcal{I}}(a)$, which implies our claim and completes the proof. Now $a\in W_{I}$ implies ${\mathcal{I}}(a)\subseteq\Phi^{+}(I)$. This, combined with the fact that ${\mathcal{I}}(w)={\mathcal{I}}(a)\sqcup a({\mathcal{I}}(d))$ by (1), yields ${\mathcal{I}}(a)\subseteq{\mathcal{I}}(w)\cap\Phi^{+}(I)$. To show the other inclusion, we first show that $d^{-1}(\Phi^{+}(I))\subseteq\Phi^{+}$. The fact that $d\in\prescript{I}{}{W}$ implies ${\mathcal{D}}_{L}(w)\cap I=\emptyset$ which in turn indicates ${\mathcal{I}}(d)\cap I=\emptyset$. Hence $d^{-1}(\alpha_{i})\subseteq\Phi^{+}$ for all $\alpha_{i}\in I$. So $d^{-1}(\Phi^{+}(I))\subseteq\Phi^{+}$ as $I$ generates $\Phi^{+}(I)$.
Now let $\alpha\in{\mathcal{I}}(w)\cap\Phi^{+}(I)$. Suppose, for contradiction, that $a^{-1}(\alpha)\in\Phi^{+}$. This, combined with $a\in W_{I}$ and $\alpha\in\Phi^{+}(I)$, means that $a^{-1}(\alpha)\in\Phi^{+}(I)$. Then, $w^{-1}(\alpha)=d^{-1}(a^{-1}(\alpha))\in\Phi^{+}$ contradicting $\alpha\in{\mathcal{I}}(w)$. So $\alpha\in{\mathcal{I}}(a)$. ∎
As an immediate corollary we have
Corollary 4.4.
Let $x\in X_{w}^{\circ}$. Then $U_{\bar{a}}\leq(B_{L_{I}})_{x}$.
We continue to work towards an understanding of the form of $(B_{L_{I}})_{x}$ by first proving a technical lemma.
@@
    $uh_{1}\notin\operatorname{Heart}_{I}(X_{w})$ for all $u\in U_{a}$ with $u\neq e$.
@@
    $tx\notin\operatorname{Heart}_{I}(X_{w})$ for all $t\in T$.
@@
(1) Since $h_{1}\in\operatorname{Heart}_{I}(X_{w})$, $h_{1}=vwB$ for some $v\in V_{d}$. By Lemma 2.3, $uv\notin V_{d}$. Thus $uvwB\in X_{w}^{\circ}\setminus\operatorname{Heart}_{I}(X_{w})$ by (5).
(2) The $T$-stability of $\operatorname{Heart}_{I}(X_{w})$, proved in Proposition  4.2, implies $tx\notin\operatorname{Heart}_{I}(X_{w})$.
(3) By Lemma 4.3, $u_{2}h_{1}=h_{1}$. Suppose, for contradiction that $u_{1}\neq e$. Then (1) implies that $u_{1}h_{1}\notin\operatorname{Heart}_{I}(X_{w})$. And thus, (2) implies $tu_{1}h_{1}\notin\operatorname{Heart}_{I}(X_{w})$. This contradicts our assumption that $tu_{1}u_{2}h_{1}=h_{2}$. Thus $u_{1}=e$ and $th_{1}=bh_{1}=h_{2}$. ∎
Finally, we are able to describe the precise form of $(B_{L_{I}})_{x}$.
@@
Let $x\in X_{w}^{\circ}$. Then $(B_{L_{I}}\cdot x)\cap\operatorname{Heart}_{I}(X_{w})\neq\emptyset$. And $(B_{L_{I}})_{x}=T_{h}\ltimes U_{\bar{a}}$ for any $h\in(B_{L_{I}}\cdot x)\cap\operatorname{Heart}_{I}(X_{w})$.
@@
Let $x\in X_{w}^{\circ}$. By (5), $x=uwB$ for $u\in U_{w}$. Thus, by Lemma 2.3, $x=u_{1}vwB$ for $u_{1}\in U_{a},v\in V_{d}$. Since $u_{1}^{-1}\in U_{a}\leq B_{L_{I}}$, we have $vwB\in B_{L_{I}}\cdot x$ and $vwB\in\operatorname{Heart}_{I}(X_{w})$ by definition. Thus $(B_{L_{I}}\cdot x)\cap\operatorname{Heart}_{I}(X_{w})\neq\emptyset$.
Let $h\in(B_{L_{I}}\cdot x)\cap\operatorname{Heart}_{I}(X_{w})$. It is immediate by Corollary 4.4 that $T_{h}\ltimes U_{\bar{a}}\leq(B_{L_{I}})_{h}$. On the other hand, let $b\in(B_{L_{I}})_{h}$ so that $bh=h$. We have $b=tu_{1}u_{2}$ with $t\in T$, $u_{1}\in U_{a}$, and $u_{2}\in U_{\bar{a}}$. Lemma 4.5(3) implies that $u_{1}=e$ and $th=h$. Hence $b\in T_{h}\ltimes U_{\bar{a}}$. We conclude that $(B_{L_{I}})_{h}=T_{h}\ltimes U_{\bar{a}}$. Our proof is complete since $(B_{L_{I}})_{x}=(B_{L_{I}})_{h}$. ∎
We now define our bijection between $B_{L_{I}}$-orbits in $X_{w}^{\circ}$ and $T$-orbits in $\operatorname{Heart}_{I}(X_{w})$.
@@
The map $\beta:\mathcal{O}_{T}(\operatorname{Heart}_{I}(X_{w}))\rightarrow\mathcal{O}_{B_{L_{I}}}(X_{w}^{\circ})$ given by $\Theta\mapsto B_{L_{I}}x$ where $x$ is any point in $\Theta$ is a codimension-preserving bijection. That is,
@@
Let $\Theta_{1},\Theta_{2}\in\mathcal{O}_{T}(\operatorname{Heart}_{I}(X_{w}))$ with $x\in\Theta_{1}$ and $y\in\Theta_{2}$. Suppose that $B_{L_{I}}\cdot x=B_{L_{I}}\cdot y$. This implies $b_{1}x=b_{2}y$ for some $b_{1},b_{2}\in B_{L_{I}}$. Thus $b_{2}^{-1}b_{1}x=y$. Lemma 4.5(3) implies there exists a $t\in T$ such that $tx=y$. Thus $\Theta_{1}=\Theta_{2}$ and $\beta$ is injective.
Let $\Xi\in\mathcal{O}_{B_{L_{I}}}(X_{w}^{\circ})$. By Lemma 4.6, there exists an $h\in\Xi\cap\operatorname{Heart}_{I}(X_{w})$. Thus $h\in\Theta$ for some $\Theta\in\mathcal{O}_{T}(\operatorname{Heart}_{I}(X_{w}))$ and so $\beta(\Theta)=\Xi$. Thus $\beta$ is surjective. Continuing with this notation, we now consider the codimension of the orbits $\Xi$ and $\Theta$ in $X_{w}^{\circ}$ and $\operatorname{Heart}_{I}(X_{w})$, respectively. One has, by [B09, Proposition 1.11], that for any algebraic group $H$ and $H$-variety $X$, the orbit $H\cdot x$, $x\in X$, is a subvariety of $X$ of dimension $\dim H-\dim H_{x}$. Thus
$\dim(\Xi)$ $=\dim(B_{L_{I}})-\dim(B_{L_{I}})_{x}$
$=\dim(B_{L_{I}})-\dim(T_{h}\ltimes U_{\bar{a}})$
$=(\ell(w_{0}(I))+\dim(T))-((\dim(T)-\dim(\Theta))+(\ell(w_{0}(I))-\ell(a)))$
@@
where the second equality is Lemma 4.6, and the third equality follows from the definition of $B_{L_{I}}$ and $U_{\bar{a}}$ as well as [B09, Proposition 1.11] applied to the $T$-orbit $\Theta$. Hence
@@
Proof of Theorem 1.4 Set $a={{}_{I}w}$ and $d=\prescript{I}{}{w}$ to match the notation we have been using throughout this section. Then [GHY24, Lemma 3.2] tells us that the set of $T$-orbits in $X^{\circ}_{d}$ for the action $\boldsymbol{\cdot}_{a^{-1}}$ is identical to the set of $T$-orbits in $X^{\circ}_{d}$ for the action $\boldsymbol{\cdot}$. Hence, Proposition 4.2 implies that the map $\mathfrak{T}:\mathcal{O}_{T}(X^{\circ}_{d})\rightarrow\mathcal{O}_{T}(\operatorname{Heart}_{I}(X_{w}))$ given by $\Theta\mapsto Tax$, where $x$ is any point in $\Theta$, is a codimension preserving bijection. The map $\mathfrak{O}$ is precisely the composition of two codimension-preserving bijections, namely $\mathfrak{O}=\beta\circ\mathfrak{T}$. Hence $\mathfrak{O}$ is itself a codimension-preserving bijection. ∎
Remark.
For the remainder of this paper, we abuse notation slightly by writing $C_{B_{L_{I}}}(X^{\circ}_{w})$ and $C_{B_{L_{I}}}(X_{w})$ for the minimum codimension of a $B_{L_{I}}$-orbit in $X^{\circ}_{w}$ and $X_{w}$, respectively. This allows for discussing these orbit codimensions even when $L_{I}$ does not act on $X^{\circ}_{w}$ or $X_{w}$.
@@
The minimal codimension of a $B_{L_{I}}$-orbit in $X^{\circ}_{w}$ equals the torus complexity of $X^{\circ}_{d}$ for the usual torus action. That is, $c_{B_{L_{I}}}(X^{\circ}_{w})=c_{T}(X^{\circ}_{d})$.

### v2

@@ Abstract.
The goal of this paper is twofold. Firstly, we provide a type-uniform formula for the torus complexity of the usual torus action on a Richardson variety, by developing the notion of algebraic dimensions of Bruhat intervals, strengthening a type $A$ result by Donten-Bury, Escobar and Portakal. In the process, we give an explicit description of the torus action on any Deodhar component as well as describe the root subgroups that comprise the component. Secondly, when a Levi subgroup in a reductive algebraic group acts on a Schubert variety, we exhibit a codimension preserving bijection between the Levi-Borel subgroup (a Borel subgroup in the Levi subgroup) orbits in the big open cell of that Schubert variety and torus orbits in the big open cell of a distinguished Schubert subvariety. This bijection has many applications including a type-uniform formula for the Levi-Borel complexity of the usual Levi-Borel subgroup action on a Schubert variety. We conclude by extending the Levi-Borel complexity results to a large class of Schubert varieties in the partial flag variety.
@@ 1.1. Group orbits in the full flag variety
Let $G$ be a complex, connected, reductive algebraic group of rank $r$. Fix a maximal torus $T$ in $G$, and a Borel subgroup $B$ of $G$ containing $T$. The homogeneous space $G/B$ is a smooth projective variety known as the full flag variety. The study of the flag variety first arose out of the need to formalize and justify the enumerative geometry of H. Schubert as laid out in Hilbert’s 15th problem. These varieties are a central theme in much of the mathematics of the 20th century and beyond, with deep and fundamental connections to algebraic geometry, Lie theory, representation theory, algebraic combinatorics, and commutative algebra.
@@ 1.3. Levi-Borel subgroup orbits in Schubert Varieties
We now turn our attention to the orbits of Levi-Borel subgroups (defined below) in a Schubert variety. Our work uncovers a deep structural relationship between the Levi-Borel orbits in the big cell of a Schubert variety and the $T$-orbits in the big cell of a distinguished Schubert subvariety. Specifically, when a Levi subgroup acts on a Schubert variety we exhibit a codimension-preserving bijection between these sets of orbits.
@@
Let $w\in W$, $I\subseteq\Delta$, and let $w={{}_{I}w}\prescript{I}{}{w}$ be the left parabolic decomposition of $w$ with respect to $I$. The map $\mathfrak{O}\!:\!\mathcal{O}_{T}(X^{\circ}_{\prescript{I}{}{w}})\!\rightarrow\!\mathcal{O}_{B_{L_{I}}}(X_{w}^{\circ})$ given by $\Theta\!\mapsto\!B_{L_{I}}\prescript{I}{}{w}x$, where $x$ is any point in $\Theta$, is a surjection. If $L_{I}$ acts on $X_{w}$, then $\mathfrak{O}$ is a codimension preserving bijection. That is,
@@
Theorem 1.4 allows us to intertwine the study of Levi-Borel orbits with the vast literature on torus orbits and varieties, including our own Corollary 1.3, and has myriad applications which we now detail. Our first application is Proposition 4.9 which provides a lower bound on the codimension of a $B_{L_{I}}$-orbit in $X_{w}$. In the case where the Levi subgroup $L_{I}$ acts on the Schubert variety $X_{w}$ this leads to a closed formula for the minimal codimension of a $B_{L_{I}}$-orbit in $X_{w}$.
@@
Our next step, with work in progress, is to use Theorem 1.4 to index the orbits of Levi-Borel subgroups in the Schubert variety itself and understand the inclusion order on their orbit closures. Throughout the years, the study of such Bruhat orders has led to much fruitful research. One related example, is that of the orbits of spherical Levi subgroups in Dynkin type $A$. These orbits are indexed by $(p,q)$-clans [MO90, Y97], and their inclusion order is described via clan combinatorics [W16].
@@
The outline of this paper is as follows. In Section 2, we introduce the necessary background, including Schubert geometry and Coxeter groups. In Section 3, we fully develop the notion of algebraic dimensions and establish Theorem 1.2. In Section 4, we prove Theorem 1.4 and Theorem 1.5 with one of our main tools being an equivariant isomorphism that allows us to use results in Section 3. Section 4 concludes with an extension of the Borel-Levi subgroup complexity results to Schubert varieties in partial flag varieties.
@@ 2.1. Combinatorics of the Weyl group
${\mathcal{I}}_{L}(w)=\{\alpha\in\Phi^{+}\>|\>w^{-1}(\alpha)\in\Phi^{-}\},\quad{\mathcal{I}}_{R}(w)=\{\alpha\in\Phi^{+}\>|\>w(\alpha)\in\Phi^{-}\},$
@@ 3. The torus complexity of Richardson varieties
In this section, we establish a formula for the complexity of the usual torus $T$ action on any Richardson variety $\mathcal{R}_{u,v}$ where $u\leq v$ in the Bruhat order.
@@ 4.1. An equivariant isomorphism
Set $U_{\bar{a}}=\prod_{\alpha\in\phi^{+}(I)\setminus{\mathcal{I}}(a)}U_{\alpha}$. If $\phi^{+}(I)\setminus{\mathcal{I}}(a)=\emptyset$, then we define $U_{\bar{a}}$ to be the trivial subgroup of $G$. Then $U_{\bar{a}}$ is a closed subgroup of $U$ normalized by $T$ and
$B_{L_{I}}=T\ltimes U_{L_{I}}=T\ltimes\prod_{\alpha\in\phi^{+}(I)}U_{\alpha}=T\ltimes(U_{\bar{a}}U_{a}).$
@@
The followings are equivalent:
  1. (1)
    $U_{\bar{a}}$ is the trivial subgroup of $G$.
  2. (2)
    $\phi^{+}(I)\setminus{\mathcal{I}}(a)=\emptyset$.
  3. (3)
    $L_{I}$ acts on $X_{w}$.
  4. (4)
    $I\subseteq\mathcal{D}_{L}(w)$.
  5. (5)
    $a=w_{0}(I)$.
@@
One direction of the equivalence between (1) and (2) is by construction and the other follows from the definition of root subgroups.
The equivalence of (3) and (4) is [BL00, Lemma 8.2.3]. If $I\subseteq{\mathcal{D}}_{L}(w)$, then ${\mathcal{I}}(w)$ contains all the positive roots in the root subsystem generated by $I$. Thus $w=w_{0}(I)d$ is a length-additive expression for some $d\in W$ by [BB05, Proposition 3.1.3]. That is, $a={{}_{I}w}=w_{0}(I)$ in the left parabolic decomposition of $w$. For the other direction, if $a=w_{0}(I)$, then $I\subseteq\mathcal{D}_{L}(w)$. Hence, (4) and (5) are equivalent.
Finally, the equivalence of (2) and (5) is immediate. ∎
Lemma 4.4.
$\operatorname{Heart}_{I}(X_{w})$ is $U_{\bar{a}}$-stable for the usual action.
Proof.
We have that
$U_{\bar{a}}\operatorname{Heart}_{I}(X_{w})$ $=U_{\bar{a}}aU_{d}a^{-1}wB=aa^{-1}U_{\bar{a}}aU_{d}dB=a\prod_{\alpha\in\phi^{+}(I)\setminus{\mathcal{I}}(a)}U_{a^{-1}(\alpha)}U_{d}dB$
$=aU_{d}dB=aU_{d}a^{-1}adB=\operatorname{Heart}_{I}(X_{w})$
where the third equality follows by [J03, Part II, 1.4(5)], and the fourth equality is (5) combined with the fact that $\alpha\notin{\mathcal{I}}(a)$ implies $U_{a^{-1}(\alpha)}\leq B$. ∎
@@
    $u_{2}h_{1}\notin\operatorname{Heart}_{I}(X_{w})$ for all $u_{2}\in U_{a}$ with $u_{2}\neq e$.
@@
    $tu_{1}x\notin\operatorname{Heart}_{I}(X_{w})$ for all $t\in T$ and $u_{1}\in U_{\bar{a}}$.
@@
(1) Since $h_{1}\in\operatorname{Heart}_{I}(X_{w})$, $h_{1}=vwB$ for some $v\in V_{d}$. By Lemma 2.3, $u_{2}v\notin V_{d}$. Thus $u_{2}vwB\in X_{w}^{\circ}\setminus\operatorname{Heart}_{I}(X_{w})$ by (5).
(2) The $T$ and $U_{\bar{a}}$ stability of $\operatorname{Heart}_{I}(X_{w})$, proved in Proposition  4.2 and Lemma 4.4 respectively, imply $tu_{1}x\notin\operatorname{Heart}_{I}(X_{w})$.
(3) Suppose, for contradiction that $u_{2}\neq e$. Then (1) implies that $u_{2}h_{1}\notin\operatorname{Heart}_{I}(X_{w})$. And thus, (2) implies $tu_{1}u_{2}h_{1}\notin\operatorname{Heart}_{I}(X_{w})$. This contradicts our assumption that $tu_{1}u_{2}h_{1}=h_{2}$. Thus $u_{2}=e$ and $tu_{1}h_{1}=bh_{1}=h_{2}$. ∎
@@
Let $x\in X_{w}^{\circ}$. Then $(B_{L_{I}}\cdot x)\cap\operatorname{Heart}_{I}(X_{w})\neq\emptyset$.
@@
Let $x\in X_{w}^{\circ}$. By (5), $x=uwB$ for $u\in U_{w}$. Thus, by Lemma 2.3, $x=u_{1}vwB$ for $u_{1}\in U_{a},v\in V_{d}$. Since $u_{1}^{-1}\in U_{a}\leq B_{L_{I}}$, we have $vwB\in B_{L_{I}}\cdot x$ and $vwB\in\operatorname{Heart}_{I}(X_{w})$ by definition. Thus $(B_{L_{I}}\cdot x)\cap\operatorname{Heart}_{I}(X_{w})\neq\emptyset$. ∎
We now define our surjection from $T$-orbits in $\operatorname{Heart}_{I}(X_{w})$ to $B_{L_{I}}$-orbits in $X_{w}^{\circ}$, which is a codimension preserving bijection in the case where $L_{I}$ acts on the Schubert variety .
@@
The map $\beta:\mathcal{O}_{T}(\operatorname{Heart}_{I}(X_{w}))\rightarrow\mathcal{O}_{B_{L_{I}}}(X_{w}^{\circ})$ given by $\Theta\mapsto B_{L_{I}}x$ where $x$ is any point in $\Theta$ is a surjection. If $L_{I}$ acts on $X_{w}$, then $\beta$ is a codimension preserving bijection. That is,
@@
Let $\Xi\in\mathcal{O}_{B_{L_{I}}}(X_{w}^{\circ})$. By Lemma 4.6, there exists an $h\in\Xi\cap\operatorname{Heart}_{I}(X_{w})$. Thus $h\in\Theta$ for some $\Theta\in\mathcal{O}_{T}(\operatorname{Heart}_{I}(X_{w}))$ and so $\beta(\Theta)=\Xi$. Thus $\beta$ is surjective.
For the remainder of this proof we assume that $L_{I}$ acts on $X_{w}$; that is, by Lemma 4.3, $U_{\bar{a}}$ is the trivial subgroup of $G$. Let $\Theta_{1},\Theta_{2}\in\mathcal{O}_{T}(\operatorname{Heart}_{I}(X_{w}))$ with $x\in\Theta_{1}$ and $y\in\Theta_{2}$. Suppose that $B_{L_{I}}\cdot x=B_{L_{I}}\cdot y$. This implies $b_{1}x=b_{2}y$ for some $b_{1},b_{2}\in B_{L_{I}}$. Thus $b_{2}^{-1}b_{1}x=y$. Lemma 4.5(3) implies there exists a $t\in T$ and $u\in U_{\bar{a}}$ such that $tux=y$. Since $U_{\bar{a}}$ is the trivial subgroup, this implies $tx=y$. Thus $\Theta_{1}=\Theta_{2}$ and $\beta$ is injective. Thus $\beta$ is a bijection.
Let $\Xi\in\mathcal{O}_{B_{L_{I}}}(X_{w}^{\circ})$. As above, there exists an $h\in\Xi\cap\operatorname{Heart}_{I}(X_{w})$ by Lemma 4.6. Thus $h\in\Theta$ for a unique $\Theta\in\mathcal{O}_{T}(\operatorname{Heart}_{I}(X_{w}))$ since $\beta$ is a bijection. We claim that
$(B_{L_{I}})_{h}=T_{h}.$
Clearly $T_{h}\subseteq(B_{L_{I}})_{h}$. For the other direction, suppose $bh=tu_{1}u_{2}h=h$ with $t\in T$, $u_{1}\in U_{\bar{a}}$, and $u_{2}\in U_{a}$. Then Lemma 4.5(3) implies $u_{2}=e$ and $tu_{1}h=h$. Since $U_{\bar{a}}$ is the trivial subgroup, $u_{1}=e$ and so $th=h$. We conclude $b=t\in T_{h}$.
One has, by [B09, Proposition 1.11], that for any algebraic group $H$ and $H$-variety $X$, the orbit $H\cdot x$, $x\in X$, is a subvariety of $X$ of dimension $\dim H-\dim H_{x}$. Thus
$\dim(\Xi)$ $=\dim(B_{L_{I}})-\dim((B_{L_{I}})_{h})$
$=\dim(B_{L_{I}})-\dim(T_{h})$
$=(\ell(w_{0}(I))+\dim(T))-((\dim(T)-\dim(\Theta))$
@@
where the fourth equality is Lemma (4.3)(4). Hence
@@
Proof of Theorem 1.4 Set $a={{}_{I}w}$ and $d=\prescript{I}{}{w}$ to match the notation we have been using throughout this section. Then [GHY24, Lemma 3.2] tells us that the set of $T$-orbits in $X^{\circ}_{d}$ for the action $\boldsymbol{\cdot}_{a^{-1}}$ is identical to the set of $T$-orbits in $X^{\circ}_{d}$ for the action $\boldsymbol{\cdot}$. Hence, Proposition 4.2 implies that the map $\mathfrak{T}:\mathcal{O}_{T}(X^{\circ}_{d})\rightarrow\mathcal{O}_{T}(\operatorname{Heart}_{I}(X_{w}))$ given by $\Theta\mapsto Tax$, where $x$ is any point in $\Theta$, is a codimension preserving bijection. The map $\mathfrak{O}$ is precisely the composition of a surjection and a bijection, namely $\mathfrak{O}=\beta\circ\mathfrak{T}$. Hence $\mathfrak{O}$ is itself a surjection. In the case where $L_{I}$ acts on $X_{w}$, $\mathfrak{O}$ is the composition of two codimension-preserving bijections, and hence is a codimension preserving bijection. ∎
@@
If $L_{I}$ acts on $X_{w}$, the minimal codimension of a $B_{L_{I}}$-orbit in $X^{\circ}_{w}$ equals the torus complexity of $X^{\circ}_{d}$ for the usual torus action. That is, $c_{L_{I}}(X^{\circ}_{w})=c_{T}(X^{\circ}_{d})$.